OpenStudy (anonymous):

Which one is a rule for the nth term of the sequence 3, 15, 75, 375...?

OpenStudy (anonymous):

By dividing consecutive terms: 15 / 3 = 5 75 /15 = 5 we find that this is a geometric series with a common ratio of 5 (k = 5). We can then plug this into the general formula for a geometric series: \[x _{n}=a _{o}\times k ^{n-1}\] where x is the term you're solving for and a is the first term. With this, we get: \[x _{n}=3 \times 5^{n-1}\]